# On the Performance of the Neyman Allocation with Small Pilots

### v3

@@ 1 Introduction
The Neyman Allocation ([31]) is a simple method for minimizing the variance of the difference-in-means estimator of ATE. In a setting without covariates, suppose that the standard deviations of the treated and control outcomes are known. The Neyman Allocation assigns more units to either treatment or control in proportion to the ratio of their standard deviations. Intuitively, the optimal experiment entails more measurements of the noisier quantity. Since the variances are not known in practice, the feasible Neyman Allocation (FNA) estimates the variances using the pilot study and then plugs the estimates into the assignment rule.
The FNA is an important part of many experiment design procedures. In the econometrics literature alone there are several notable works. [24, ] propose to estimate the variance of outcome and control groups conditional on covariate value, implementing the FNA conditional on covariates. In a similar vein, [34, ] employs tree-based techniques to stratify units based on their covariates. Units are then assigned to treatment and control based on the FNA conditional on strata. [18, ] proposes local randomization to select representative units for participation and treatment in experiments. In what [18, ] terms the “fully efficient” case, treatment proportion conditional on the randomization group is chosen by the FNA. Despite their differences, the above papers study their proposals in asymptotic frameworks that take the pilot size to infinity. Their analyses, appropriate for large pilots, essentially assume that population parameters are arbitrarily well-estimated from the pilots alone. In practice, pilots are often conducted for logistical reasons and may be small. In such settings, accurately estimating the relevant variances may be difficult. Indeed, all of the above authors caution against the use of their methods when pilot sizes are small.
@@
The above findings suggest that researchers should be cautious when designing experiments using the FNA with small pilots. However, even when pilots are large, methods which condition on many covariates may end up estimating the FNA using a small conditional sample. Furthermore, if researchers believe that units exhibit cluster-dependence – a common assumption in empirical work – the number of “effective” observations may be smaller still, impeding the estimation of the FNA.
@@
However, our results do not imply that the FNA should never be used. Instead, researchers can consider alternative procedures with better performance under homoskedasticity. We explore three candidate solutions via simulations. One based on testing for homoskedasticity, and two that continuously regularize the FNA towards the balanced allocation. Our simulations suggest that the procedure we term “exponential regularization” seems to offer the best trade-offs in that performance under heteroskedasticity is the least compromised in exchange for better performance under homoskedasticity. Alternatively, researchers might consider optimizing other aspects of the design, such as how strata are formed, or in the selection of covariates for stratification. [5, ] and [19, ] discuss methods to robustify these approaches for small pilots.
We note that our paper specifically addresses the use of the FNA in reducing the asymptotic variance of the difference-in-means estimator in two-wave experiments. It does not speak to papers which take the treatment assignment probability as given, such as [5, ].
@@
The challenges that small pilots pose for the FNA was noted as early as [33, ], which is concerned with small pilots and small main wave experiment. Given the relative intractability of the set-up, the author concludes using simulations that the FNA performs well. Authors such as [24, 34, 5, ] have noted the limitations of their method in a small pilot setting. In particular, [18, 19, ] highlights that the FNA leads to estimators with residual variation from variance estimation when pilots are small. [19, ] further shows that estimating stratification variables using small pilots may be counterproductive. In a network setting, [35, ] characterizes the regret of the designed experiment as a function of pilot size, providing a rule for choosing pilot size. Our paper builds on this body of work by setting up balanced randomization as the natural alternative to the FNA when pilots are small, focusing on the relative efficiency of the two methods.
@@
Our paper stands in contrast to [6, ], which argues that the FNA can perform well even when pilots are small. However, their theoretical results rely on pilot sizes going to infinity, which we consider less suitable for addressing the small pilot case. Their simulation results, which considers only normally distributed outcomes, may understate the challenges arising from high kurtosis.
@@
Our paper is closely related to papers pointing out a similar issue in the design of sequential experiments. [30, ] argue by simulation that treatment assignment rules based on estimated outcomes can do worse than non-adaptive rules due to estimation noise. Theoretical analysis is provided in [25, ], in an asymptotic framework that does not nest ours. Our paper is also related to those studying small sample problems in experiments. [15, ] is concerned with the problems small strata pose for inference. [9, ] considers the effectiveness of various randomization strategies in achieving balance when a single-wave experiment is small. Finally, we note that there is a large literature discussing the large-pilot properties of the Neyman Allocation for alternative criteria such as power (e.g. [8], [3]), minimax optimality (e.g. [4]) or ethical considerations (e.g. Chapter 8 of [26]). These criteria fall outside the scope of this present paper.
@@ 2 Framework
þ Potential outcomes $(Y(0),Y(1))\sim F$ have finite second moments. The vector of means is $\mu$ and the covariance matrix is $\Sigma$ where
@@
þ Potential outcomes in the pilot, denoted $\left\{\widetilde{Y}_{i}(0),\widetilde{Y}_{i}(1)\right\}_{i=1}^{m}$, consist of $m$ i.i.d. draws from the distribution of the random vector $(Y(0),Y(1))^{\prime}$. Treatment is randomly assigned in the pilot so that denoting assignments by $\left\{\widetilde{A}_{i}\right\}_{i=1}^{m}$,
@@
þ Potential outcomes in the main wave, denoted by $\left\{Y_{i}(0),Y_{i}(1)\right\}_{i=1}^{n}$, are $n$ i.i.d. draws from the distribution of the random vector $(Y(0),Y(1))^{\prime}$ and are independent to potential outcomes and treatment assignments in the pilot. That is,
@@
    $\widehat{\theta}^{\text{rand}}_{p}=\frac{\frac{1}{n}\sum_{i=1}^{n}Y_{p,i}A^{\text{rand}}_{p,i}}{\frac{1}{n}\sum_{i=1}^{n}A^{\text{rand}}_{p,i}}-\frac{\frac{1}{n}\sum_{i=1}^{n}Y_{p,i}\left(1-A^{\text{rand}}_{p,i}\right)}{\frac{1}{n}\sum_{i=1}^{n}\left(1-A^{\text{rand}}_{p,i}\right)}\leavevmode\nobreak\ .$
  • Block/Complete Randomization: For a given $p\in(0,1)$, let $n_{1}=\lfloor np\rfloor$. Randomly assign exactly $n_{1}$ units to treatment so that all ($n\choose n_{1}$ assignments are equally likely. In other words, let $A_{p}=(A_{p,1},\dots,A_{p,n})^{\prime}$ denote the random vector of treatment assignments. Then, under block/complete randomization, the distribution of $A_{p}^{(n)}$ satisfies
    $P(A_{p}=\mathbf{a})=1/{n\choose n_{1}}\text{ for all }\mathbf{a}\in\{0,1\}^{n}\text{ such that }\sum_{i=1}^{n}a_{i}=n_{1}.$
@@
    $\widehat{\theta}_{p}=\frac{\frac{1}{n}\sum_{i=1}^{n}Y_{p,i}A_{p,i}}{\frac{1}{n}\sum_{i=1}^{n}A_{p,i}}-\frac{\frac{1}{n}\sum_{i=1}^{n}Y_{p,i}\left(1-A_{p,i}\right)}{\frac{1}{n}\sum_{i=1}^{n}\left(1-A_{p,i}\right)}\leavevmode\nobreak\ .$
@@
  • The Infeasible Neyman Allocation: For any given $p\in(0,1)$, elementary arguments show that
@@
    $p_{\ast}=\frac{\sigma(1)}{\sigma(1)+\sigma(0)}\leavevmode\nobreak\ .$
@@
    $\begin{split}Y_{p_{\ast},i}=&\ Y_{i}(0)\left(1-A_{p_{\ast},i}\right)+Y_{i}(1)A_{p_{\ast},i}\\ \widehat{\theta}_{p_{\ast}}=&\ \frac{\frac{1}{n}\sum_{i=1}^{n}Y_{p_{\ast},i}A_{p_{\ast},i}}{\frac{1}{n}\sum_{i=1}^{n}A_{p_{\ast},i}}-\frac{\frac{1}{n}\sum_{i=1}^{n}Y_{p_{\ast},i}\left(1-A_{p_{\ast},i}\right)}{\frac{1}{n}\sum_{i=1}^{n}\left(1-A_{p_{\ast},i}\right)}\leavevmode\nobreak\ .\end{split}$ (2.1)
@@
    $\widetilde{\sigma}_{m}^{2}(a)=\ \frac{1}{m_{a}-1}\sum_{i=1}^{m}\left(\widetilde{Y}_{i}\mathbb{I}\left\{\widetilde{A}_{i}=a\right\}-\frac{1}{m_{a}}\sum_{i=1}^{m}\widetilde{Y}_{i}\mathbb{I}\left\{\widetilde{A}_{i}=a\right\}\right)^{2}\leavevmode\nobreak\ ,$
    where $m_{a}=\ \sum_{i=1}^{m}\mathbb{I}\left\{\widetilde{A}_{i}=a\right\}$. The Feasible Neyman Allocation is
    $\widetilde{p}=\begin{cases}\frac{\widetilde{\sigma}_{m}(1)}{\widetilde{\sigma}_{m}(1)+\widetilde{\sigma}_{m}(0)}&\text{if }\widetilde{\sigma}_{m}(1),\widetilde{\sigma}_{m}(0)>0\leavevmode\nobreak\ ,\\ \frac{1}{2}&\text{otherwise}\leavevmode\nobreak\ .\end{cases}$ (2.2)
@@
    $\begin{split}Y_{\widetilde{p},i}=&\ Y_{i}(0)\left(1-A_{\widetilde{p},i}\right)+Y_{i}(1)A_{\widetilde{p},i}\\ \widehat{\theta}_{\widetilde{p}}=&\ \frac{\frac{1}{n}\sum_{i=1}^{n}Y_{\widetilde{p},i}A_{\widetilde{p},i}}{\frac{1}{n}\sum_{i=1}^{n}A_{\widetilde{p},i}}-\frac{\frac{1}{n}\sum_{i=1}^{n}Y_{\widetilde{p},i}\left(1-A_{\widetilde{p},i}\right)}{\frac{1}{n}\sum_{i=1}^{n}\left(1-A_{\widetilde{p},i}\right)}\leavevmode\nobreak\ .\end{split}$ (2.3)
@@
We can also define INA and FNA with simple random assignment instead of block/complete randomization. However, the latter is generally considered to be preferable to simple random assignment (see e.g. [28]) and is the standard in fields such as Development Economics (see examples in [11]).
@@ 3 Toy Example
$\begin{pmatrix}Y(1)\\ Y(0)\end{pmatrix}\sim{N}(\mu,\Sigma)\leavevmode\nobreak\ .$
@@
proportion of the units in the main to treatment. Suppose for simplicity again that we assign the first $n\widetilde{p}$ units to treatment and the remainder to control (rounding $n\widetilde{p}$ if necessary). Then,
$\widehat{\theta}_{\widetilde{p}}$ $=\left(\frac{1}{n\widetilde{p}}\sum_{i=1}^{n\widetilde{p}}Y_{i}(1)\right)-\left(\frac{1}{n-n\widetilde{p}}\sum_{i=n\widetilde{p}+1}^{n}Y_{i}(0)\right)\leavevmode\nobreak\ .$
@@
$\left[\frac{1}{\sqrt{n\widetilde{p}}}\sum_{i=1}^{n\widetilde{p}}Y_{i}(1)\;\bigg{\lvert}\;\widetilde{p}\right]\sim{N}\left(0,\sigma^{2}(1)\right)\quad,\quad\left[\frac{1}{\sqrt{n-n\widetilde{p}}}\sum_{i=n\widetilde{p}+1}^{n}Y_{i}(0)\;\bigg{\lvert}\;\widetilde{p}\right]\sim{N}\left(0,\sigma^{2}(0)\right)\leavevmode\nobreak\ .$
@@
$\left[\sqrt{n}\left(\widehat{\theta}_{\widetilde{p}}-\theta\right)\;\bigg{\lvert}\;\widetilde{p}\right]\sim{N}\left(0\;,\;\frac{\sigma^{2}(1)}{\widetilde{p}}+\frac{\sigma^{2}(0)}{1-\widetilde{p}}\right)\leavevmode\nobreak\ .$
@@
$\text{Var}\left[\sqrt{n}\left(\widehat{\theta}_{\widetilde{p}}-\theta\right)\right]=\mathbb{E}\left[\frac{\sigma^{2}(1)}{\widetilde{p}}+\frac{\sigma^{2}(0)}{1-\widetilde{p}}\right]\leavevmode\nobreak\ .$ (3.1)
@@
$\text{Var}\left[\sqrt{n}\left(\widehat{\theta}_{p}-\theta\right)\right]=\frac{\sigma^{2}(1)}{p}+\frac{\sigma^{2}(0)}{1-p}\leavevmode\nobreak\ .$
@@
$\mathbb{E}\left[\frac{\sigma^{2}(1)}{\widetilde{p}}+\frac{\sigma^{2}(0)}{1-\widetilde{p}}\right]$ $=\mathbb{E}\left[\left(1+\frac{1}{Z_{m}}\frac{\sigma(0)}{\sigma(1)}\right)\sigma^{2}(1)+\left(1+Z_{m}\frac{\sigma(1)}{\sigma(0)}\right)\sigma^{2}(0)\right]\leavevmode\nobreak\ ,$
@@
$\frac{\sigma(1)}{\sigma(0)}\in C_{m}:=\left[\;\mathbb{E}\left[Z_{m}\right]-\sqrt{\mathbb{E}\left[Z_{m}\right]^{2}-1}\;,\;\mathbb{E}\left[Z_{m}\right]+\sqrt{\mathbb{E}\left[Z_{m}\right]^{2}-1}\;\right]\leavevmode\nobreak\ .$ (3.2)
@@
$\mathbb{E}[Z_{m}]^{2}=\mathbb{E}[Z_{m}]\mathbb{E}\left[\frac{1}{Z_{m}}\right]>1\leavevmode\nobreak\ .$
@@
$\left(\mathbb{E}\left[Z_{m}\right]+\sqrt{\mathbb{E}\left[Z_{m}\right]^{2}-1}\right)\left(\mathbb{E}\left[Z_{m}\right]-\sqrt{\mathbb{E}\left[Z_{m}\right]^{2}-1}\right)=1\leavevmode\nobreak\ .$
@@
$C_{m}=\left[\frac{1}{c_{m}},c_{m}\right]\leavevmode\nobreak\ ,$
@@
Figure 1: $C_{m}$ when $Y(1)\sim{N}(\mu(1),\sigma^{2}(1))$ and $Y(0)\sim{N}(\mu(0),\sigma^{2}(0))$. Monte Carlo integration using 10,000 draws.
@@
As our toy model shows, it is the estimation of $\widetilde{\sigma}(1)$ and $\widetilde{\sigma}(0)$ that causes problems when $m$ is small. It is therefore intuitive that $C_{m}$ will be wide when the distributions are fat tailed, that is, when kurtosis is high. As such, we consider the following parametrization: $Y(1)\sim\mu(1)+\sigma(1)\cdot\text{Pareto}(l,s)$ and $Y(0)\sim\mu(0)+\sigma(0)\cdot\text{Pareto}(l,s)$. Here, $l$ and $s$ are the location and scale parameters respectively. Figure 2 plots $C_{m}$ for $l=1$ and $s\in\{3,4,5\}$. Indeed, we see that the bands are much wider than in the previous models. For $s=3$, even when $m=50$, $C_{m}=[0.50,2.15]$, which, given the examples in Section 5 appear to be fairly extreme amounts of heteroskedasticity. Finally, we note that, $C_{m}$ decreases in width as we move from $s=3$ to $s=5$.
Figure 2: $C_{m}$ when $Y(1)\sim\mu(1)+\sigma(1)\cdot\text{Pareto}(1,s)$, $Y(0)\sim\mu(0)+\sigma(0)\cdot\text{Pareto}(1,s)$ and $s\in\{3,4,5\}$. Monte Carlo integration using 10,000 draws.

### v4

@@ 1 Introduction
The Neyman Allocation (Neyman 1934) is a simple method for minimizing the variance of the difference-in-means estimator of ATE. In a setting without covariates, suppose that the standard deviations of the treated and control outcomes are known. The Neyman Allocation assigns more units to either treatment or control in proportion to the ratio of their standard deviations. Intuitively, the optimal experiment entails more measurements of the noisier quantity. Since the variances are not known in practice, the feasible Neyman Allocation (FNA) estimates the variances using the pilot study and then plugs the estimates into the assignment rule.
The FNA is an important part of many experiment design procedures. In the econometrics literature alone there are several notable works. Hahn et al. (2011) propose to estimate the variance of outcome and control groups conditional on covariate value, implementing the FNA conditional on covariates. In a similar vein, Tabord-Meehan (2021) employs tree-based techniques to stratify units based on their covariates. Units are then assigned to treatment and control based on the FNA conditional on strata. Cytrynbaum (2021) proposes local randomization to select representative units for participation and treatment in experiments. In what Cytrynbaum (2021) terms the “fully efficient" case, treatment proportion conditional on the randomization group is chosen by the FNA. Despite their differences, the above papers study their proposals in asymptotic frameworks that take the pilot size to infinity. Their analyses, appropriate for large pilots, essentially assume that population parameters are arbitrarily well-estimated from the pilots alone. In practice, pilots are often conducted for logistical reasons and may be small. In such settings, accurately estimating the relevant variances may be difficult. Indeed, all of the above authors caution against the use of their methods when pilot sizes are small.
@@
The above findings suggest that researchers should be cautious when designing experiments using the FNA with small pilots. However, even when pilots are large, methods which condition on many covariates may end up estimating the FNA using a small conditional sample. Furthermore, if researchers believe that units exhibit cluster-dependence – a common assumption in empirical work – the number of “effective" observations may be smaller still, impeding the estimation of the FNA.
@@
However, our results do not imply that the FNA should never be used. Instead, researchers may consider alternative procedures with better performance under homoskedasticity. We explore three candidate solutions via simulations. One is based on testing for homoskedasticity, and the other two continuously regularize the FNA towards the balanced allocation. Our simulations suggest that the procedure we term “exponential regularization" offers the best trade-offs in that performance under heteroskedasticity is the least compromised in exchange for better performance under homoskedasticity. Alternatively, researchers might consider optimizing other aspects of the design, such as how strata are formed, or in the selection of covariates for stratification. Bai (2022) and Cytrynbaum (2023) discuss methods to robustify these approaches for small pilots.
@@
The challenges that small pilots pose for the FNA were noted as early as Sukhatme (1935), which is concerned with both small pilots and small main wave experiments. Given the relative intractability of the set-up, the author concludes using simulations that the FNA performs well. Authors such as Hahn et al. (2011); Tabord-Meehan (2021); Bai (2022) have noted the limitations of their method in a small pilot setting. In particular, Cytrynbaum (2021, 2023) highlights that the FNA leads to estimators with residual variation from variance estimation when pilots are small. Cytrynbaum (2023) further shows that estimating stratification variables using small pilots may be counterproductive. In a network setting, Viviano (2022) characterizes the regret of the designed experiment as a function of pilot size, providing a rule for choosing pilot size. Our paper builds on this body of work by setting up balanced randomization as the natural alternative to the FNA when pilots are small, focusing on the relative efficiency of the two methods. Our analysis specifically addresses the use of the FNA in reducing the asymptotic variance of the difference-in-means estimator in two-wave experiments. It does not speak to papers which take the treatment assignment probability as given, such as Bai (2022).
@@
Our work stands in contrast to Blackwell et al. (2022), who argue that the FNA can perform well even when pilots are small. However, their theoretical results rely on pilot sizes going to infinity, which we consider less suitable for addressing the small pilot case. Their simulations, which consider only normally distributed outcomes, may understate the challenges arising from high kurtosis.
@@
Our paper is closely related to papers pointing out a similar issue in the design of sequential experiments. Melfi and Page (1998) argue by simulation that treatment assignment rules based on estimated outcomes can do worse than non-adaptive rules due to estimation noise. Theoretical analysis is provided in Hu and Rosenberger (2003), in an asymptotic framework that does not nest ours. Our paper is also related to those studying small sample problems in experiments. de Chaisemartin and Ramirez-Cuellar (2019) are concerned with the problems small strata pose for inference. Bruhn and McKenzie (2009) consider the effectiveness of various randomization strategies in achieving balance when a single-wave experiment is small. Finally, we note that there is a large literature discussing the large-pilot properties of the Neyman Allocation for alternative criteria such as power (e.g. Brittain and Schlesselman 1982, Azriel et al. 2012), minimax optimality (e.g. Bai 2021) or ethical considerations (e.g. Chapter 8 of Hu and Rosenberger 2006). These criteria fall outside the scope of this present paper.
@@ 2 Framework
Potential outcomes $(Y(0),Y(1))\sim F$ have finite second moments. The vector of means is $\mu$ and the covariance matrix is $\Sigma$ where
@@
Potential outcomes in the pilot, denoted $\left\{\widetilde{Y}_{i}(0),\widetilde{Y}_{i}(1)\right\}_{i=1}^{m}$, consist of $m$ i.i.d. draws from the distribution of the random vector $(Y(0),Y(1))^{\prime}$. Treatment is randomly assigned in the pilot so that denoting assignments by $\left\{\widetilde{A}_{i}\right\}_{i=1}^{m}$,
@@
Potential outcomes in the main wave, denoted by $\left\{Y_{i}(0),Y_{i}(1)\right\}_{i=1}^{n}$, are $n$ i.i.d. draws from the distribution of the random vector $(Y(0),Y(1))^{\prime}$ and are independent to potential outcomes and treatment assignments in the pilot. That is,
@@
    $\widehat{\theta}^{\text{rand}}_{p}=\frac{\frac{1}{n}\sum_{i=1}^{n}Y_{p,i}A^{\text{rand}}_{p,i}}{\frac{1}{n}\sum_{i=1}^{n}A^{\text{rand}}_{p,i}}-\frac{\frac{1}{n}\sum_{i=1}^{n}Y_{p,i}\left(1-A^{\text{rand}}_{p,i}\right)}{\frac{1}{n}\sum_{i=1}^{n}\left(1-A^{\text{rand}}_{p,i}\right)}~{}.$
  • Block/Complete Randomization: For a given $p\in(0,1)$, let $n_{1}=\lfloor np\rfloor$. Randomly assign exactly $n_{1}$ units to treatment so that all $\binom{n}{n_{1}}$ assignments are equally likely. In other words, let $A_{p}=(A_{p,1},\dots,A_{p,n})^{\prime}$ denote the random vector of treatment assignments. Then, under block/complete randomization, the distribution of $A_{p}$ satisfies
    $P(A_{p}=\mathbf{a})=1/{\binom{n}{n_{1}}}\text{ for all }\mathbf{a}\in\{0,1\}^{n}\text{ such that }\sum_{i=1}^{n}a_{i}=n_{1}.$
@@
    $\widehat{\theta}_{p}=\frac{\frac{1}{n}\sum_{i=1}^{n}Y_{p,i}A_{p,i}}{\frac{1}{n}\sum_{i=1}^{n}A_{p,i}}-\frac{\frac{1}{n}\sum_{i=1}^{n}Y_{p,i}\left(1-A_{p,i}\right)}{\frac{1}{n}\sum_{i=1}^{n}\left(1-A_{p,i}\right)}~{}.$
@@
  • The Infeasible Neyman Allocation: For any given $p\in(0,1)$, it can be shown that
@@
    $p_{\ast}=\frac{\sigma(1)}{\sigma(1)+\sigma(0)}~{}.$
@@
    $\begin{split}Y_{p_{\ast},i}=&\ Y_{i}(0)\left(1-A_{p_{\ast},i}\right)+Y_{i}(1)A_{p_{\ast},i}\\ \widehat{\theta}_{p_{\ast}}=&\ \frac{\frac{1}{n}\sum_{i=1}^{n}Y_{p_{\ast},i}A_{p_{\ast},i}}{\frac{1}{n}\sum_{i=1}^{n}A_{p_{\ast},i}}-\frac{\frac{1}{n}\sum_{i=1}^{n}Y_{p_{\ast},i}\left(1-A_{p_{\ast},i}\right)}{\frac{1}{n}\sum_{i=1}^{n}\left(1-A_{p_{\ast},i}\right)}~{}.\end{split}$
@@
    $\widetilde{\sigma}_{m}^{2}(a)=$ $\,\frac{1}{m_{a}-1}\sum_{i=1}^{m}\left(\widetilde{Y}_{i}\mathbb{I}\left\{\widetilde{A}_{i}=a\right\}-\frac{1}{m_{a}}\sum_{i=1}^{m}\widetilde{Y}_{i}\mathbb{I}\left\{\widetilde{A}_{i}=a\right\}\right)^{2}~{},$
    where $\text{where }\widetilde{Y}_{i}=$ $\,\widetilde{Y}_{i}(1)\widetilde{A}_{i}+\widetilde{Y}_{i}(0)\left(1-\widetilde{A}_{i}\right)~{},$
    and $\text{and }m_{a}=$ $\,\sum_{i=1}^{m}\mathbb{I}\left\{\widetilde{A}_{i}=a\right\}~{}.$
    The Feasible Neyman Allocation is
    $\widetilde{p}=\begin{cases}\frac{\widetilde{\sigma}_{m}(1)}{\widetilde{\sigma}_{m}(1)+\widetilde{\sigma}_{m}(0)}&\text{if }\widetilde{\sigma}_{m}(1),\widetilde{\sigma}_{m}(0)>0~{},\\ \frac{1}{2}&\text{otherwise}~{}.\end{cases}$ (2.1)
@@
    $\begin{split}Y_{\widetilde{p},i}=&\ Y_{i}(0)\left(1-A_{\widetilde{p},i}\right)+Y_{i}(1)A_{\widetilde{p},i}~{},\\ \widehat{\theta}_{\widetilde{p}}=&\ \frac{\frac{1}{n}\sum_{i=1}^{n}Y_{\widetilde{p},i}A_{\widetilde{p},i}}{\frac{1}{n}\sum_{i=1}^{n}A_{\widetilde{p},i}}-\frac{\frac{1}{n}\sum_{i=1}^{n}Y_{\widetilde{p},i}\left(1-A_{\widetilde{p},i}\right)}{\frac{1}{n}\sum_{i=1}^{n}\left(1-A_{\widetilde{p},i}\right)}~{}.\end{split}$
@@
We can also define INA and FNA with simple random assignment instead of block/complete randomization. However, the latter is generally considered to be preferable to simple random assignment (see e.g. Lachin (1988)) and is the standard in fields such as Development Economics (see examples in Bugni et al. (2018)).
@@ 3 Toy Example
$\begin{pmatrix}Y(1)\\ Y(0)\end{pmatrix}\sim{N}(\mu,\Sigma)~{}.$
@@
proportion of the units in the main wave to treatment. Suppose for simplicity again these are the first $n\widetilde{p}$ units (rounding $n\widetilde{p}$ if necessary). Then,
$\widehat{\theta}_{\widetilde{p}}=\left(\frac{1}{n\widetilde{p}}\sum_{i=1}^{n\widetilde{p}}Y_{i}(1)\right)-\left(\frac{1}{n-n\widetilde{p}}\sum_{i=n\widetilde{p}+1}^{n}Y_{i}(0)\right)~{}.$
@@
$\left[\frac{1}{\sqrt{n\widetilde{p}}}\sum_{i=1}^{n\widetilde{p}}Y_{i}(1)\;\bigg{\lvert}\;\widetilde{p}\right]\sim$ $\,N\left(0,\sigma^{2}(1)\right)~{},$
$\left[\frac{1}{\sqrt{n-n\widetilde{p}}}\sum_{i=n\widetilde{p}+1}^{n}Y_{i}(0)\;\bigg{\lvert}\;\widetilde{p}\right]\sim$ $\,N\left(0,\sigma^{2}(0)\right)~{}.$
@@
$\left[\sqrt{n}\left(\widehat{\theta}_{\widetilde{p}}-\theta\right)\;\bigg{\lvert}\;\widetilde{p}\right]\sim{N}\left(0\;,\;\frac{\sigma^{2}(1)}{\widetilde{p}}+\frac{\sigma^{2}(0)}{1-\widetilde{p}}\right)~{}.$
@@
$\text{Var}\left[\sqrt{n}\left(\widehat{\theta}_{\widetilde{p}}-\theta\right)\right]=\mathbb{E}\left[\frac{\sigma^{2}(1)}{\widetilde{p}}+\frac{\sigma^{2}(0)}{1-\widetilde{p}}\right]~{}.$ (3.1)
@@
$\text{Var}\left[\sqrt{n}\left(\widehat{\theta}_{p}-\theta\right)\right]=\frac{\sigma^{2}(1)}{p}+\frac{\sigma^{2}(0)}{1-p}~{}.$
@@
$\mathbb{E}\left[\frac{\sigma^{2}(1)}{\widetilde{p}}+\frac{\sigma^{2}(0)}{1-\widetilde{p}}\right]$ $=\mathbb{E}\left[\left(1+\frac{1}{Z_{m}}\frac{\sigma(0)}{\sigma(1)}\right)\sigma^{2}(1)+\left(1+Z_{m}\frac{\sigma(1)}{\sigma(0)}\right)\sigma^{2}(0)\right]~{},$
@@
$\frac{\sigma(1)}{\sigma(0)}\in C_{m}:=\left[\;\mathbb{E}\left[Z_{m}\right]-\sqrt{\mathbb{E}\left[Z_{m}\right]^{2}-1}\;,\;\mathbb{E}\left[Z_{m}\right]+\sqrt{\mathbb{E}\left[Z_{m}\right]^{2}-1}\;\right]~{}.$ (3.2)
@@
$\mathbb{E}[Z_{m}]^{2}=\mathbb{E}[Z_{m}]\mathbb{E}\left[\frac{1}{Z_{m}}\right]>1~{}.$
@@
$\left(\mathbb{E}\left[Z_{m}\right]+\sqrt{\mathbb{E}\left[Z_{m}\right]^{2}-1}\right)\left(\mathbb{E}\left[Z_{m}\right]-\sqrt{\mathbb{E}\left[Z_{m}\right]^{2}-1}\right)=1~{}.$
@@
$C_{m}=\left[\frac{1}{c_{m}},c_{m}\right]~{},$
@@
Figure 1: $C_{m}$ when $Y(a)\sim\mu(a)+\sigma(a)\cdot N(0,1)$ and $Y(a)\sim\mu(a)+\sigma(a)\cdot\chi^{2}_{1}$ for $a\in\{0,1\}$. Monte Carlo integration using 10,000 draws.
@@
As our toy model shows, estimation of $\widetilde{\sigma}(1)$ and $\widetilde{\sigma}(0)$ causes problems when $m$ is small. It is therefore intuitive that $C_{m}$ will be wide when the distributions are fat tailed, that is, when kurtosis is high. As such, we consider the following parametrization: $Y(1)\sim\mu(1)+\sigma(1)\cdot\text{Pareto}(l,s)$ and $Y(0)\sim\mu(0)+\sigma(0)\cdot\text{Pareto}(l,s)$. Here, $l$ and $s$ are the location and scale parameters respectively. Figure 2 plots $C_{m}$ for $l=1$ and $s\in\{3,4,5\}$. Indeed, we see that the bands are much wider than in the previous models. For $s=3$, even when $m=50$, $C_{m}=[0.50,2.15]$. Given the examples in Section 5, this appears to be fairly extreme amounts of heteroskedasticity. Finally, we note that, $C_{m}$ decreases in width as we move from $s=3$ to $s=5$.
Figure 2: $C_{m}$ when $Y(a)\sim\mu(a)+\sigma(a)\cdot\text{Pareto}(1,s)$, for $a\in\{0,1\}$ and $s\in\{3,4,5\}$. Monte Carlo integration using 10,000 draws.